Consider the following data for a dependent variable y and two independent variables, x1 and x2.

x1 x2 y
29 13 94
46 10 109
24 17 112
50 17 178
40 6 94
52 19 176
75 8 171
36 13 118
60 13 143
76 17 212

Round your all answers to two decimal places. Enter negative values as negative numbers, if necessary.

a. Develop an estimated regression equation relating y to x1 .

y = __ + __ x1

Predict y if x1 = 45

y=___

  

b. Develop an estimated regression equation relating y to x2.

      

Predict y if x2 = 25

y = __.

  

c. Develop an estimated regression equation relating y to x1 and x2.

y = ___ + ___x1 + ___x2   

Predict y if x1=45 and x2=25 .

y = ___
Answer #1

( a )

Using Excel we get output as

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.812209997
R Square 0.659685079
Adjusted R Square 0.617145714
Standard Error 25.55354199
Observations 10
ANOVA
df SS MS F Significance F
Regression 1 10126.23193 10126.23 15.50764 0.00430814
Residual 8 5223.868066 652.9835
Total 9 15350.1
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 48.21953469 24.83565128 1.941545 0.088136 -9.0515798 105.490649
x1 1.895091502 0.481235166 3.937974 0.004308 0.78536122 3.00482178

from the above output

y = 48.22 + 1.90 x1

if x1 = 45

y = 48.22 + 1.90 ( 45 )

= 133.72

( b )

Using Excel we get output as

Regression Statistics
Multiple R 0.502203567
R Square 0.252208422
Adjusted R Square 0.158734475
Standard Error 37.87920851
Observations 10
ANOVA
df SS MS F Significance F
Regression 1 3871.424503 3871.425 2.698168 0.13908867
Residual 8 11478.6755 1434.834
Total 9 15350.1
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept 76.49006623 40.8842998 1.870891 0.098272 -17.789298 170.769431
x2 4.82781457 2.939111618 1.64261 0.139089 -1.949789 11.6054181

from the above output

y = 76.49 + 4.83 x2

if x2 = 25

y = 76.49 + 4.83 ( 25 )

= 197.24

( c )

Using Excel we get output as

Regression Statistics
Multiple R 0.972662657
R Square 0.946072645
Adjusted R Square 0.930664829
Standard Error 10.87454629
Observations 10
ANOVA
df SS MS F Significance F
Regression 2 14522.3097 7261.155 61.40213 3.6419E-05
Residual 7 827.7902994 118.2558
Total 9 15350.1
Coefficients Standard Error t Stat P-value Lower 95% Upper 95%
Intercept -22.69958904 15.71624081 -1.44434 0.191871 -59.862593 14.4634151
x1 1.94512482 0.204958423 9.490339 3.02E-05 1.46047516 2.42977448
x2 5.148684046 0.844451487 6.097075 0.000492 3.15187358 7.14549451

from the above output

y = -22.70 + 1.95 x1 + 5.15 x2

= -22 70 + 1.95 ( 45 ) + 5.15 ( 25 )

= 193.8